### v5

@@ 3 Bipartite dependent rounding algorithm
We consider here a variant of a rounding algorithm from [8], which was in turn inspired by an earlier “fair sharing” algorithm of [4]. For motivation, consider the following natural rounding procedure: each edge $e$ draws an Exponential random $Z_{e}$ variable with rate $x_{e}$. Then, for each right-node $v$, we set $X_{f}=1$ for the edge $f=\operatorname{argmin}_{e\in\Gamma(u)}X_{e}$. It is easy to see that this is equivalent to independent rounding, and in particular we have $\Pr(X_{e}=1)=x_{e}$.
@@
1 for each left-node $u$ do
2       Call $\vec{Z}^{(u)}\leftarrow\textsc{CorrelatedExponential}(\vec{\rho}^{(u)})$ for vector $\langle\rho^{u}(v)=\rho_{(u,v)}:v\in\Gamma(u)\rangle$
3Combine all vectors $\vec{Z}^{(u)}$ into a single vector $Z\in\mathbb{R}_{\geq 0}^{|E|}$ with $Z_{(u,v)}=Z^{(u)}_{v}$ for all $u,v$.
@@ 6       Set $X_{e}=1$ and $X_{e^{\prime}}=0$ for all other neighbors $e^{\prime}\in\Gamma(v)\setminus\{e\}$
Suppose that $\vec{\rho}\in[0,1]^{|E|}$ satisfies $\sum_{e\in\Gamma(u)}\rho_{e}=1$ for all left-nodes $u$. Then DepRound can be implemented in polynomial time with the following properties:
@@
Define $L^{\prime}$ to be the set of edges outside $L$ which share a right-node with an edge in $L$, that, is, the set of edges of the form $f=(u,v)\notin L$ where $(u^{\prime},v)\in L$. Let $W,W^{\prime}$ denote the set of left-nodes of edges of $L,L^{\prime}$ respectively. We claim that $W,W^{\prime}$ are disjoint. For, suppose $u\in W\cap W^{\prime}$. So there edges $(u,v)\in L,(u,v^{\prime})\in L^{\prime}$; by definition of $L^{\prime}$, there must be a corresponding edge $(u^{\prime},v^{\prime})\in L$. Since $G$ is a simple graph, necessarily $u\neq u^{\prime}$. Also $v\neq v^{\prime}$ since edges in $L$ have distinct right-nodes. Thus $(u,v)\sim(u^{\prime},v^{\prime})$, contradicting that $L$ is a stable set.
@@
$\Pr(\bigwedge_{e\in L}X_{e}=1)=\Pr(\bigwedge_{(u,v)\in L}Z_{(u,v)}/x_{(u,v)}=\min_{f\in\Gamma(v)}Z_{f}/x_{f})$
@@
$\Pr(\bigwedge_{(u,v)\in L}Z_{(u,v)}/x_{u,v}=\min_{f\in\Gamma(v)}Z_{f}/x_{f})\leq\prod_{(u,v)\in L}\Pr(Z_{(u,v)}/x_{(u,v)}=\min_{f\in\Gamma(v)}Z_{f}/x_{f})$
@@
Consider a stable edge set $L\subseteq E$, and let $\mathcal{E}$ denote the event that $X_{e}=1$ for all $e\in L$. By iterated expectations with respect to random variable $Z_{L}$ and Lemma 11, we have
$\Pr(\mathcal{E})={\mathbb{E}}_{Z_{L}}[\Pr(\mathcal{E}\mid Z_{L})]\leq{\mathbb{E}}\bigl{[}\prod_{e\in L}\mathrm{e}^{(1-1/x_{e})Z_{e}}\bigr{]}$
@@
[…]
@@
For any edges $e_{1}=(u,v_{1}),e_{2}=(u,v_{2})$ with the same left-node $u$ and $\rho_{e_{1}},\rho_{e_{2}}\in[0,1)$, there holds
${\mathbb{E}}[X_{1}X_{2}]\leq x_{1}x_{2}\cdot\Bigl{(}1-\frac{((1-\rho_{1})^{1-1/x_{1}}-1)((1-\rho_{2})^{1-1/x_{2}}-1)}{(1-\rho_{1})^{1-1/x_{1}}(1-\rho_{2})^{1-1/x_{2}}+\rho_{1}+\rho_{2}-1}\Bigr{)}$
where we write $X_{i}=X_{e_{i}},x_{i}=x_{e_{i}},\rho_{i}=\rho_{e_{i}}$ for $i=1,2$ for brevity.
@@
Let $Z_{i}=Z_{e_{i}}$ for $i=1,2$. By Lemma 11 applied to stable-set $L=\{e_{1},e_{2}\}$ we have:
@@
[…]
@@
There is not any clear optimal choice of vector $\vec{\rho}$. However, there is one attractive option which leads to a particularly strong and algebraically simple formulation of property (A4):
@@
Suppose that for edges $e_{1}=(u,v_{1}),e_{2}=(u,v_{2})$ with the same left-node $u$ we set $\rho_{e_{i}}=x_{e_{i}}t$ for some parameter $t\geq 4/3$. Then
${\mathbb{E}}[X_{1}X_{2}]\leq x_{1}x_{2}\Bigl{(}1-\frac{\mathrm{e}^{t}-1}{\mathrm{e}^{t}+1}+0.57(x_{1}+x_{2})\max\{0,1/t-0.45\}\Bigr{)}$
where we write $X_{i}=X_{e_{i}},x_{i}=x_{e_{i}}$ for brevity.
@@
Suppose that for edges $e_{1}=(u,v_{1}),e_{2}=(u,v_{2})$ with the same left-node $u$ we set $\rho_{e_{i}}=1-\mathrm{e}^{-x_{e_{i}}t}\leq x_{e_{i}}t$ for some parameter $t>0$. Then we have:
${\mathbb{E}}[X_{1}X_{2}]\leq x_{1}x_{2}\cdot\frac{(\mathrm{e}^{t}-1)(\mathrm{e}^{x_{1}t}+\mathrm{e}^{x_{2}t})}{\mathrm{e}^{2t}-\mathrm{e}^{x_{1}t}-\mathrm{e}^{x_{2}t}+\mathrm{e}^{(x_{1}+x_{2})t}}\leq x_{1}x_{2}\cdot\frac{\mathrm{e}^{x_{1}t}+\mathrm{e}^{x_{2}t}}{1+\mathrm{e}^{t}}$
where we write $X_{i}=X_{e_{i}},x_{i}=x_{e_{i}}$ for brevity.
@@
The result is obvious is either $x_{1}$ or $x_{2}$ is equal to zero, and so we may suppose that $\rho_{e_{1}},\rho_{e_{2}}\in(0,1)$ strictly. Let $\eta=\mathrm{e}^{t}$. Now, for a given vertex $u$ and edges $e_{1}=(u,v_{1}),e_{2}=(u,v_{2})$ with the same left-node $u$, Theorem 13 with some algebraic simplifications gives
${\mathbb{E}}[X_{1}X_{2}]\leq x_{1}x_{2}\cdot\frac{(1-\rho_{1})^{1-1/x_{1}}+(1-\rho_{2})^{1-1/x_{2}}+\rho_{1}+\rho_{2}-2}{(1-\rho_{1})^{1-1/x_{1}}(1-\rho_{2})^{1-1/x_{2}}+\rho_{1}+\rho_{2}-1}=x_{1}x_{2}\frac{(\eta-1)(\eta^{x_{1}}+\eta^{x_{2}})}{\eta^{2}-\eta^{x_{1}}-\eta^{x_{2}}+\eta^{x_{1}+x_{2}}}$
@@ 4 Scheduling to minimize completion time
1 Define parameters $\pi=3.9,\theta=0.555,\tau=0.604$.
@@
8            Initialize cluster index $\ell=1$ and set $\mathcal{C}^{(i)}_{k,1}=\emptyset$.
9            Sort the jobs in $\mathcal{P}^{(i)}_{k}$ in non-increasing order of Smith ratio as $j_{1},j_{2},\dots,j_{s}$.
@@
16                        Update $\ell\leftarrow\ell+1$ and initialize the new cluster $\mathcal{C}^{(i)}_{k,\ell}=\emptyset$
@@
So each machine has a single “open” cluster for each processing-time quantization class at a time. If, after adding the job to the open cluster, the cluster size becomes at least $\theta$, then we close it and open a new one. The correlation parameter $\rho^{(i)}_{j}$ is usually chosen to be proportional to $x^{(i)}_{j}$; the one exception is that the final job that closes out a cluster may need to choose a smaller parameter $\rho^{(i)}_{j}$. We say a job $j$ is truncated on a machine $i$ if $\tilde{\rho}_{j}^{(i)}<x^{(i)}_{j}$.
Note that, in our rounding algorithm, we only use the diagonal terms $x^{(i)}_{j}$ of the SDP relaxation; the “cross-terms” $x^{(i)}_{j,j^{\prime}}$ appear only for the analysis. We also remark that, unlike prior algorithms, there is no special handling for “large” jobs (jobs with large mass $x^{(i)}_{j}$). For sake of notational convenience, we will suppose throughout that $x^{(i)}_{j}\in(0,1)$ strictly for all $i,j$.
@@
${\mathbb{E}}[X^{(i)}_{j}X^{(i)}_{j^{\prime}}]\leq(1-\phi^{(i)}_{j,j^{\prime}})\cdot x^{(i)}_{j}x^{(i)}_{j^{\prime}}\quad\text{ for $\phi^{(i)}_{j,j^{\prime}}:=\frac{((1-\rho^{(i)}_{j})^{1-1/x^{(i)}_{j}}-1)((1-\rho^{(i)}_{j^{\prime}})^{1-1/x^{(i)}_{j^{\prime}}}-1)}{(1-\rho^{(i)}_{j})^{1-1/x^{(i)}_{j}}(1-\rho^{(i)}_{j^{\prime}})^{1-1/x^{(i)}_{j^{\prime}}}+\rho^{(i)}_{j}+\rho^{(i)}_{j^{\prime}}-1}$}$
@@
Apply Corollary 2 to the graph with left-vertices $\mathcal{M}$ and right-vertices $\mathcal{J}$ and blocks $\mathcal{C}^{(i)}_{k,\ell}$, with $x_{e}=x^{(i)}_{j},\rho_{e}=\rho^{(i)}_{j},X^{(i)}_{j}=X_{e}$ for each edge $e=(i,j)$. Note that $\sum_{e\in\Gamma(j)}x_{e}=\sum_{e\in\mathcal{C}^{(i)}_{k,\ell}}\rho_{e}=1$.
By Property (A1), every job is assigned to exactly one machine. By Property (A2), we have ${\mathbb{E}}[X^{(i)}_{j}]=x^{(i)}_{j}$. By Property (A3’), every pair of jobs $j,j^{\prime}$ on a machine $i$ has ${\mathbb{E}}[X^{(i)}_{j}X^{(i)}_{j^{\prime}}]\leq x^{(i)}_{j}x^{(i)}_{j^{\prime}}$. By Property (A4’), every pair of jobs $j,j^{\prime}$ in a cluster $\mathcal{C}^{(i)}_{k,\ell}$ has ${\mathbb{E}}[X^{(i)}_{j}X^{(i)}_{j^{\prime}}]\leq(1-\phi^{(i)}_{j,j^{\prime}})x^{(i)}_{j}x^{(i)}_{j^{\prime}}$. ∎
@@
Note that $H^{(i)}_{j}\in[1,\pi]$ for all $j$. The key property we exploit is that each random value $\log_{\pi}H^{(i)}_{j}$ is uniformly distributed in $[0,1]$. The following observation is what we use for calculations:

### v8

@@ 3 Bipartite dependent rounding algorithm
We consider here a variant of a rounding algorithm from [9], which was in turn inspired by an earlier “fair sharing” algorithm of [5]. For motivation, consider the following natural rounding procedure: each edge $e$ draws an Exponential random $Z_{e}$ variable with rate $x_{e}$. Then, for each right-node $v$, we set $X_{f}=1$ for the edge $f=\operatorname{argmin}_{e\in\Gamma(v)}X_{e}$. It is easy to see that this is equivalent to independent rounding, and in particular we have $\Pr(X_{e}=1)=x_{e}$.
@@
1for each left-node $u$ do
2   Call $\vec{Z}^{(u)}\leftarrow\textsc{CorrelatedExponential}(\vec{\rho}^{(u)})$ for vector $(\rho^{u}(v)=\rho_{(u,v)}:v\in\Gamma(u))$
3Combine all vectors $\vec{Z}^{(u)}$ into a single vector $Z\in\mathbb{R}_{\geq 0}^{E}$ with $Z_{(u,v)}=Z^{(u)}_{v}$ for all $u,v$.
@@ 6   Set $X_{e}=1$ and $X_{e^{\prime}}=0$ for all other neighbors $e^{\prime}\in\Gamma(v)\setminus\{e\}$
Suppose that $\vec{\rho}\in[0,1]^{E}$ satisfies $\sum_{e\in\Gamma(u)}\rho_{e}=1$ for all left-nodes $u$. Then DepRound can be implemented in polynomial time with the following properties:
@@
Define $L^{\prime}$ to be the set of edges outside $L$ which share a right-node with an edge in $L$, that, is, the set of edges of the form $f=(u,v)\notin L$ where $(u^{\prime},v)\in L$. Let $W,W^{\prime}$ denote the set of left-nodes of edges of $L,L^{\prime}$ respectively. We claim that $W,W^{\prime}$ are disjoint. For, suppose $u\in W\cap W^{\prime}$. So there edges $(u,v)\in L,(u,v^{\prime})\in L^{\prime}$; by definition of $L^{\prime}$, there must be a corresponding edge $(u^{\prime},v^{\prime})\in L$. Since $G$ is a simple graph, necessarily $u\neq u^{\prime}$. Also $v\neq v^{\prime}$ since edges in $L$ have distinct right-nodes. This contradicts that $L$ is a stable set.
@@
$\Pr(\bigwedge_{e\in L}X_{e}=1)=\Pr\Bigl(\bigwedge_{(u,v)\in L}Z_{(u,v)}/x_{(u,v)}=\min_{f\in\Gamma(v)}Z_{f}/x_{f}\Bigr)$
@@
$\Pr\Bigl(\bigwedge_{(u,v)\in L}Z_{(u,v)}/x_{u,v}=\min_{f\in\Gamma(v)}Z_{f}/x_{f}\Bigr)\leq\prod_{(u,v)\in L}\Pr\bigl(Z_{(u,v)}/x_{(u,v)}=\min_{f\in\Gamma(v)}Z_{f}/x_{f}\bigr)$
@@
Consider a stable edge set $L\subseteq E$. By iterated expectations with respect to random variable $Z_{L}$ and Lemma 11, we have
${\mathbb{E}}\Bigl[\prod_{e\in L}X_{e}\Bigr]={\mathbb{E}}_{Z_{L}}\Bigl[{\mathbb{E}}\bigl[\prod_{e\in L}X_{e}\mid Z_{L}\bigr]\Bigr]\leq{\mathbb{E}}\Bigl[\prod_{e\in L}\mathrm{e}^{(1-1/x_{e})Z_{e}}\Bigr]$
@@
Here, for an edge $e$, we have ${\mathbb{E}}[\mathrm{e}^{(1-1/x_{e})Z_{e}}]=\int_{z=0}^{\infty}{\mathrm{e}}^{-z}\cdot\mathrm{e}^{(1-1/x_{e})z}\ \mathrm{d}z=x_{e}.$
@@
For any edges $e_{1}=(u,v_{1}),e_{2}=(u,v_{2})$ with the same left-node $u$, there holds ${\mathbb{E}}[X_{e_{1}}X_{e_{2}}]\leq x_{e_{1}}x_{e_{2}}\cdot(1-\Phi(x_{e_{1}},x_{e_{2}};\rho_{e_{1}},\rho_{e_{2}}))$.
@@
Let us write $X_{i}=X_{e_{i}},x_{i}=x_{e_{i}},\rho_{i}=\rho_{e_{i}},Z_{i}=Z_{e_{i}}$ for $i=1,2$. By Lemma 11 applied to stable-set $L=\{e_{1},e_{2}\}$ we have:
@@
Observe that values $Z_{1},Z_{2}$ are simultaneously generated by $\textsc{CorrelatedExponential}(\vec{\rho}^{(u)})$. By applying Lemma 8 with $q_{1}=1-1/x_{1},q_{2}=1-1/x_{2}$, we have
@@
As we have mentioned, there is no single optimal choice for the vector $\vec{\rho}$. One attractive option is to simply set $\vec{\rho}$ proportional to $\vec{p}$, which leads to the following simple version of property (A4):
@@
For $t\geq 4/3$ and $x_{1},x_{2}\in(0,1)$, there holds
$\Phi(x_{1},x_{2};tx_{1},tx_{2})\geq 1-\frac{\mathrm{e}^{t}-1}{\mathrm{e}^{t}+1}+0.57(x_{1}+x_{2})\max\{0,1/t-0.45\}$
@@
For $t>0$ and $x_{1},x_{2}\in(0,1)$, there holds
$\Phi(x_{1},x_{2};1-\mathrm{e}^{-x_{1}t},1-\mathrm{e}^{-x_{2}t})=1-\frac{(\mathrm{e}^{t}-1)(\mathrm{e}^{x_{1}t}+\mathrm{e}^{x_{2}t})}{\mathrm{e}^{2t}-\mathrm{e}^{x_{1}t}-\mathrm{e}^{x_{2}t}+\mathrm{e}^{(x_{1}+x_{2})t}}\geq 1-\frac{\mathrm{e}^{x_{1}t}+\mathrm{e}^{x_{2}t}}{1+\mathrm{e}^{t}}$
@@
Let $\rho_{i}=1-\mathrm{e}^{x_{i}t}$ for $i=1,2$, and let $\eta=\mathrm{e}^{t}$. With some algebraic simplifications we get
$\Phi(x_{1},x_{2};\rho_{1},\rho_{2})=1-\frac{(\mathrm{e}^{t}-1)(\mathrm{e}^{x_{1}t}+\mathrm{e}^{x_{2}t})}{\mathrm{e}^{2t}-\mathrm{e}^{x_{1}t}-\mathrm{e}^{x_{2}t}+\mathrm{e}^{(x_{1}+x_{2})t}}=1-\frac{(\eta-1)(\eta^{x_{1}}+\eta^{x_{2}})}{\eta^{2}-\eta^{x_{1}}-\eta^{x_{2}}+\eta^{x_{1}+x_{2}}}$
@@ 4 Scheduling to minimize completion time
1Define parameters $\pi=3.9,\theta=0.555,\tau=0.604$.
@@
8     Initialize cluster index $\ell=1$ and set $\mathcal{C}^{(i)}_{k,1}$$\emptyset$.
9     Sort the jobs in $\mathcal{P}^{(i)}_{k}$ with $x^{(i)}_{j}>0$ in non-increasing order of Smith ratio as $j_{1},j_{2},\dots,j_{s}$.
@@
16           Update $\ell\leftarrow\ell+1$ and initialize the new cluster $\mathcal{C}^{(i)}_{k,\ell}$$\emptyset$
@@
So each machine has a single “open” cluster for each processing-time quantization class at a time. If, after adding the job to the open cluster, the cluster size becomes at least $\theta$, then we close it and open a new one. The correlation parameter $\rho^{(i)}_{j}$ is usually chosen to be proportional to $x^{(i)}_{j}$; the one exception is that the final job that closes out a cluster may need to choose a smaller parameter $\rho^{(i)}_{j}$. We say a job $j$ is truncated on a machine $i$ if $\tilde{\rho}_{j}^{(i)}<x^{(i)}_{j}$; note that, in this case, every other job $j^{\prime}$ in the cluster has $\rho_{j^{\prime}}^{(i)}=x^{(i)}_{j}/\tau$.
[…]
@@
${\mathbb{E}}[X^{(i)}_{j}X^{(i)}_{j^{\prime}}]\leq(1-\phi^{(i)}_{j,j^{\prime}})\cdot x^{(i)}_{j}x^{(i)}_{j^{\prime}}\quad\text{ for $\phi^{(i)}_{j,j^{\prime}}:=\Phi(x^{(i)}_{j},x^{(i)}_{j^{\prime}};\rho^{(i)}_{j},\rho^{(i)}_{j^{\prime}})$}$
@@
Apply Theorem 1 to the graph with left-nodes $\mathcal{C}^{(i)}_{k,\ell}$ and right-nodes $\mathcal{J}$; for each machine $i$ and job $j$ where $j\in\mathcal{C}^{(i)}_{k,\ell}$, it has an edge $e=(\mathcal{C}^{(i)}_{k,\ell},j)$ with $x_{e}=x^{(i)}_{j},\rho_{e}=\rho^{(i)}_{j},X^{(i)}_{j}=X_{e}$. Note that $\sum_{e\in\Gamma(j)}x_{e}=\sum_{e\in\mathcal{C}^{(i)}_{k,\ell}}\rho_{e}=1$.
By Property (A1), every job is assigned to exactly one machine. By Property (A2), we have ${\mathbb{E}}[X^{(i)}_{j}]=x^{(i)}_{j}$. For any pair of jobs $j,j^{\prime}$ on a machine $i$, note that the corresponding edge-set $\{(\mathcal{C}^{(i)}_{k,\ell},j),(\mathcal{C}^{(i)}_{k^{\prime},\ell^{\prime}},j^{\prime})\}$ is stable. So Property (A3) gives ${\mathbb{E}}[X^{(i)}_{j}X^{(i)}_{j^{\prime}}]\leq x^{(i)}_{j}x^{(i)}_{j^{\prime}}$. Also, by Property (A4), every pair of jobs $j,j^{\prime}$ in a cluster $\mathcal{C}^{(i)}_{k,\ell}$ has ${\mathbb{E}}[X^{(i)}_{j}X^{(i)}_{j^{\prime}}]\leq(1-\phi^{(i)}_{j,j^{\prime}})x^{(i)}_{j}x^{(i)}_{j^{\prime}}$. ∎
@@
Note that $H^{(i)}_{j}\in[1,\pi]$ for all $j$. The key property we exploit is that each random value $\log_{\pi}H^{(i)}_{j}$ is uniformly distributed in $[0,1]$. This leads to the following formula: